Calculate ∆G° for the reaction occurring in the following cell :

Al (s) | Al³⁺ 1(M) || Cu²⁺ 1(M) Cu (s)

Standard electrode potentials :

E°= −1.66 V and E°= + 0.34 V
Al³⁺/Al Cu²⁺/Cu
1
Expert's answer
2019-09-23T03:48:40-0400

E°=0.34-(-1.66)=2 V

∆G°=-nFE°/1000=-6*96485*2=-1157 kJ/mol